38. A consulting group recently conducted a global survey of product teams with the goal of better understanding the dynamics of product team performance and uncovering the practices that make these teams successful. One of the survey findings was that 34​% of organizations have a coherent business strategy that they stick to and effectively communicate. Suppose another study is conducted to check the validity of this​ result, with the goal of proving that the percentage is less than 34​%.

b. A sample of 150 organizations is​ selected, and results indicate that 51 organizations have a coherent business strategy that they stick to and effectively communicate. Use the​ p-value approach to determine at the 0.01 level of significance whether there is evidence that the percentage is less than 34%​

Calculate the test statistic.

Calculate p value.

State the conclusion.
Answer #1

Answer:

Given,

Null hypothesis Ho : p = 0.34

Alternative hypothesis Ha : p < 0.34

sample proportion p^ = x/n

= 51/150

= 0.34

test statistic z = (p^ - p)/sqrt(pq/n)

substitute values

= (0.34 - 0.34)/sqrt(0.34(1-0.34)/150)

z = 0

P value = P(z < 0)

= 0.5

Here we observe that, p value > alpha(0.01), so we fail to reject Ho.

So there is no sufficient evidence.